# GSQA: An End-to-End Model for Generative Spoken Question Answering

### v2

@@ Abstract
In recent advancements in spoken question answering (SQA), end-to-end models have made significant strides. However, previous research has primarily focused on extractive span selection. While this extractive-based approach is effective when answers are present directly within the input, it falls short in addressing abstractive questions, where answers are not directly extracted but inferred from the given information. To bridge this gap, we introduce the first end-to-end Generative Spoken Question Answering (GSQA) model that empowers the system to engage in abstractive reasoning. The challenge in training our GSQA model lies in the absence of a spoken abstractive QA dataset. We propose using text models for initialization and leveraging the extractive QA dataset to transfer knowledge from the text generative model to the spoken generative model. Experimental results indicate that our model surpasses the previous extractive model by 3% on extractive QA datasets. Furthermore, the GSQA model has only been fine-tuned on the spoken extractive QA dataset. Despite not having seen any spoken abstractive QA data, it can still closely match the performance of the cascade model. In conclusion, our GSQA model shows the potential to generalize to a broad spectrum of questions, thus further expanding the SQA capabilities of abstractive QA. Our code is available at https://voidful.github.io/GSQA
@@ 1 Introduction
Recently, end-to-end, textless approaches (e.g., DUAL[5]) have emerged as a solution to the cascade method’s limitations. These works [5, 6, 7] rely on the encoded representations of acoustic models, like HuBERT[8], and apply K-means clustering to transform speech representations into discrete units. These units, encapsulating both questions and passages, are fed into the model that predicts the span of the answer in the passage using start and end positions. By adopting this textless approach, DUAL effectively sidesteps the error propagation from ASR systems [9, 10, 11]. It improves the model robustness in noisy speech. However, there’s a notable limitation to the DUAL model: it is not able to handle abstractive QA scenarios, as it is designed to only output start and end positions within the input spoken passage.
@@ 3.2 Speech Quantization
HuBERT [8] is used for efficient speech representation, converting speech signals into discrete units through mask segment prediction training. This involves clustering each 20ms frame into one of K categories using k-means clustering, where the centroid ID of each cluster represents the frame’s discrete unit. We utilize the HuBERT-base model with a 100-cluster configuration, following Lee et al.’s methods [6], to achieve this conversion. Additionally, we employ run-length encoding to condense consecutive identical units into a single unit, thus reducing the length of input unit sequences.
@@ 3.3 Text Question Answering Pretraining
We select LongT5 [19] for this purpose. LongT5, an extension of the T5 encoder, incorporates global-local attention mechanisms to efficiently manage longer inputs. It combines the attention efficiency of ETC with PEGASUS’s summarization-focused pre-training, offering improved performance. We further pretrain LongT5 on textual QA datasets, which would be denoted as LongT5-TQA.
@@ 3.4 GSQA: Generative Spoken QA Model
Fig.2 illustrates our GSQA training procedure. As mentioned in Sec.3.3, we pretrain the LongT5 on Text Question Answering (TQA) tasks to produce LongT5-TQA, which is capable of text-level extractive and abstractive QA. This model is then fine-tuned on the Extractive Spoken QA dataset, NMSQA [5]. To unify the dataset format in abstractive QA form, we modify the labels in NMSQA from time spans to the answers’ waveform. Here, the given spoken passages, questions, and answers would go through a quantization model, and would be converted into discrete units. Since LongT5-TQA does not recognize discrete units, we introduced them as new tokens. The embeddings of units are initialized by random sampling from other text token embeddings, ensuring proper tokenization and a seamless transition from text to speech. By these modifications, LongT5-TQA would be fine-tuned to predict answers in discrete units instead of text, where we denote the fine-tuned model as Unit-LongT5-TQA. Note that if the model’s denotation is not appended with TQA, it means that we don’t pretrain the model on TQA tasks. Last, to achieve speech-to-speech, we employ a HiFi-GAN vocoder [20] to convert discrete units back to speech. This vocoder includes a duration prediction module for deduplicated unit synthesis.
@@ 4.1 Dataset
[…]
@@ 4.2 Cascaded Pipeline
The conventional approach for SQA tasks typically follows a cascaded pipeline, as illustrated in the middle of Fig. 1. The passages and questions are first transcribed by an ASR system, followed by a text-based language model that generates text-based answers, and a TTS system converts the text-based answers into speech answers. To be more specific, we use Whisper-medium.en [13] as the ASR system, the fine-tuned LongT5-TQA model as the text-based language model, and the Azure TTS service as the TTS system. We record the experiment result of the cascaded pipeline in the column of Cascade Model in Table 1. The total number of parameters of the cascaded model is shown in Table 2, and the WERs of the ASR system in given datasets are listed in Table 3.
@@ 6 Conclusion
In conclusion, we introduce the first end-to-end Generative Spoken Question Answering (GSQA) model, bridging the gap between spoken extractive and abstractive QA tasks. By combining the advantages of cascaded and textless models, our GSQA model demonstrates superior performance on extractive QA datasets, outperforming the previous E2E extractive SQA models. Furthermore, in the challenging abstractive zero-shot domain, our model exhibits competitive capabilities. We also highlighted the importance of pretraining the language model on textual data to enhance the unit-to-unit model’s capability of semantic understanding. Our future research would concentrate on improving the model’s performance with human narrations and exploring training strategies to enhance generalization ability across diverse SQA tasks.

### v4

@@ Abstract
In recent advancements in spoken question answering (SQA), end-to-end models have made significant strides. However, previous research has primarily focused on extractive span selection. While this extractive-based approach is effective when answers are present directly within the input, it falls short in addressing abstractive questions, where answers are not directly extracted but inferred from the given information. To bridge this gap, we introduce the first end-to-end Generative Spoken Question Answering (GSQA) model that empowers the system to engage in abstractive reasoning. The challenge in training our GSQA model lies in the absence of a spoken abstractive QA dataset. We propose using text models for initialization and leveraging the extractive QA dataset to transfer knowledge from the text generative model to the spoken generative model. Experimental results indicate that our model surpasses the previous extractive model by 3% on extractive QA datasets. Furthermore, the GSQA model has only been fine-tuned on the spoken extractive QA dataset. Despite not having seen any spoken abstractive QA data, it can still closely match the performance of the cascade model. In conclusion, our GSQA model shows the potential to generalize to a broad spectrum of questions, thus further expanding the SQA capabilities of abstractive QA.
@@ 1 Introduction
Recently, end-to-end, textless approaches (e.g., DUAL[5]) have emerged as a solution to the cascade method’s limitations. These works [5, 6, 7] rely on the encoded representations of acoustic models, like HuBERT[8], and apply K-means clustering to transform speech representations into discrete units. These units, encapsulating both questions and passages, are fed into the model that predicts the span of the answer in the passage using start and end positions. By adopting this textless approach, DUAL effectively sidesteps the error propagation from ASR systems [9, 10, 11]. It improves the model robustness in noisy speech. However, there is a notable limitation to the DUAL model: it is not able to handle abstractive QA scenarios, as it is designed to only output start and end positions within the input spoken passage.
@@ 3.2 Speech Quantization
HuBERT [8] is used for efficient speech representation, converting speech signals into discrete units through mask segment prediction training. This involves clustering each 20ms frame into one of K categories using k-means clustering, where the centroid ID of each cluster represents the frame’s discrete unit. To achieve this conversion, we utilize the HuBERT-base model with a 100-cluster configuration, following Lee et al.’s methods [6]. Additionally, we employ run-length encoding to condense consecutive identical units into a single unit, thus reducing the length of input unit sequences.
@@ 3.3 Text Question Answering Pretraining
We select LongT5 [19] for this purpose. LongT5, an extension of the T5 encoder, incorporates global-local attention mechanisms to efficiently manage longer inputs. It combines the attention efficiency of ETC with PEGASUS’s summarization-focused pre-training, offering improved performance. We further pretrain LongT5 on textual QA datasets, denoted as LongT5-TQA.
@@ 3.4 GSQA: Generative Spoken QA Model
Fig.2 illustrates our GSQA training procedure. As mentioned in Sec.3.3, we pretrain the LongT5 on Text Question Answering (TQA) tasks to produce LongT5-TQA, which is capable of text-level extractive and abstractive QA. This model is then fine-tuned on the Extractive Spoken QA dataset, NMSQA [5]. To unify the dataset format in abstractive QA form, we modify the labels in NMSQA from time spans to the answers’ waveform. Here, the given spoken passages, questions, and answers would go through a quantization model and would be converted into discrete units. Since LongT5-TQA does not recognize discrete units, we introduced them as new tokens. The embeddings of units are initialized by random sampling from other text token embeddings, ensuring proper tokenization and a seamless transition from text to speech. By these modifications, LongT5-TQA would be fine-tuned to predict answers in discrete units instead of text, where we denote the fine-tuned model as Unit-LongT5-TQA. Note that if the model’s denotation is not appended with TQA, it means that we don’t pretrain the model on TQA tasks. Last, to achieve speech-to-speech, we employ a HiFi-GAN vocoder [20] to convert discrete units back to speech. This vocoder includes a duration prediction module for deduplicated unit synthesis.
@@ 4.1 Dataset
Method Speech-to-(text/unit) model QA Model Total Cascade model Whisper-medium.en (764M) LongT5-TQA (261M) 1025M DUAL HuBERT-large-128-layer1-22 (304M) LongFormer (148M) 452M Unit-LongT5-TQA HuBERT-base-100-layer1-6 (51M) LongT5-TQA (261M) 312M
@@ 4.2 Cascaded Pipeline
The conventional approach for SQA tasks typically follows a cascaded pipeline, as illustrated in the middle of Fig. 1. The passages and questions are first transcribed by an ASR system, followed by a text-based language model that generates text-based answers, and a TTS system converts the text-based answers into speech answers. To be more specific, we use Whisper-medium.en [13] as the ASR system, the fine-tuned LongT5-TQA model as the text-based language model, and the Azure TTS service as the TTS system. We record the cascaded pipeline’s experiment result in the Cascade Model column in Table 1. The total number of parameters of the cascaded model is shown in Table 2, and the WERs of the ASR system in given datasets are listed in Table 3.
@@ 6 Conclusion
In conclusion, we introduce the first end-to-end Generative Spoken Question Answering (GSQA) model, bridging the gap between spoken extractive and abstractive QA tasks. Combining the advantages of cascaded and textless models, our GSQA model demonstrates superior performance on extractive QA datasets, outperforming the previous E2E extractive SQA models. Furthermore, in the challenging abstractive zero-shot domain, our model exhibits competitive capabilities. We also highlighted the importance of pretraining the language model on textual data to enhance the unit-to-unit model’s capability of semantic understanding. Our future research would concentrate on improving the model’s performance with human narrations and exploring training strategies to enhance generalization ability across diverse SQA tasks.
7 Acknowledgement
We thank the National Center for High-performance Computing (NCHC) of National Applied Research Laboratories (NARLabs) in Taiwan for providing computational and storage resources.